OpenStudy (ammarah):

Write the first six terms of the sequence. 1. an =n^2+6

hartnn (hartnn):

you know how to get the first term ?

OpenStudy (ammarah):

plug in 1?

hartnn (hartnn):

yes! put n= 1 in an =n^2+6 you will get a1 which is 1st term

hartnn (hartnn):

in exactly same manner, if you plug in n=2, you get a2 which is 2nd term plug in n=3, get 3rd term....and so on, till you plug in n=6 to get sixth term try it

OpenStudy (ammarah):

ok i got it.

hartnn (hartnn):

good! :)

OpenStudy (ammarah):

f(n)=-4/3n this means the same thing right?

hartnn (hartnn):

you want first few terms of f(n) too ? to get first term, put n= 1, to get f(1) and so on

OpenStudy (ammarah):

so if i put 2 in it would be -4/6

hartnn (hartnn):

yes, and that will be 2nd term = f(2)

OpenStudy (ammarah):

ok imma tagg u in another question

hartnn (hartnn):

ok :)